OpenStudy (anonymous):

What is the difference between a ratio and a proportion? A. A proportion is a comparison of two quantities, and a ratio is a an equation where proportions are set equal. B. A ratio is a comparison of two quantities, and a proportion is an equation where ratios are set equal. C. They are the same concept D. They are exact opposites

OpenStudy (anonymous):

Ratio: A ratio is a comparison of two numbers. We generally separate the two numbers in the ratio with a colon (:). Suppose we want to write the ratio of 8 and 12. We can write this as 8:12 or as a fraction 8/12, and we say the ratio is eight to twelve. Example: Jeannine has a bag with 3 videocassettes, 4 marbles, 7 books, and 1 orange. Proportion: A proportion is an equation with a ratio on each side. It is a statement that two ratios are equal. 3/4 = 6/8 is an example of a proportion. When one of the four numbers in a proportion is unknown, cross products may be used to find the unknown number. This is called solving the proportion. Question marks or letters are frequently used in place of the unknown number. Example: Solve for n: 1/2 = n/4. Using cross products we see that 2 × n = 1 × 4 =4, so 2 × n = 4. Dividing both sides by 2, n = 4 ÷ 2 so that n = 2. Hope that helps :)

OpenStudy (anonymous):

so it would be B

OpenStudy (anonymous):

:) yes

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

welcome :)